Thesaurus of key words used in the annual subject indexes (valid from January 1997)

The list is common to Astronomy and Astrophysics, The Astrophysical Journal and Monthly Notices of the Royal Astronomical Society. In order to ease the search, the key words are subdivided into broad categories.

No more than six codes all together should be listed for a paper as this is the limit fixed by the computer program.

The key words in boldface listed under the code numbers 07.03.2, 07.16.2, 08.09.2, 08.16.5, 08.19.5, 09.09.1, 09.16.2, 10.07.3, 10.15.2, 11.02.2, 11.09.1, 11.17.4 are intended for use with specific astronomical objects; each contains the word “individual”. The corresponding code numbers should never be used alone, but always in combination with the most common names for the astronomical objects in question. For example, if a paper discusses three individual galaxies, these should be coded on the title page of the manuscript in the following manner:

11.09.1 Arp 220; 11.09.1 M 51; 11.09.1 NGC 4472

Note that each object (in the example the three galaxies) counts as one code within the allowed limit of six codes.

The parts of the key words in italics are for reference only and should be omitted when the key words are entered on the manuscript.

General

01.02.1 Book reviews 01.05.1 Editorials notices 01.05.2 Errata, addenda 01.05.3 Extraterrestrial intelligence 01.08.1 History and philosophy of astronomy 01.13.1 Miscellaneous 01.15.1 Obituaries, biographies

Physical data and processes

02.01.1 Acceleration of particles 02.01.2 Accretion, accretion disks 02.01.3 Atomic data 02.01.4 Atomic processes 02.02.1 Black hole physics 02.03.1 Chaos 02.03.2 Conduction 02.03.3 Convection 02.03.4 Cosmic strings 02.04.1 Dense matter 02.04.2 Diffusion 02.05.1 Elementary particles 02.05.2 Equation of state 02.07.1 Gravitation 02.07.2 Gravitational waves 02.08.1 Hydrodynamics 02.09.1 Instabilities 02.12.1 Line: formation 02.12.2 Line: identification 02.12.3 Line: profiles 02.13.1 Magnetic fields 02.13.2 Magnetohydrodynamics (MHD) 02.13.3 Masers 02.13.4 Molecular data 02.13.5 Molecular processes 02.14.1 Nuclear reactions, nucleosynthesis, abundances 02.16.1 Plasmas 02.16.2 Polarization 02.18.5 Radiation mechanisms: non-thermal 02.18.6 Radiation mechanisms: thermal 02.18.7 Radiative transfer 02.18.8 Relativity 02.19.2 Scattering 02.19.1 Shock waves 02.20.1 Turbulence 02.23.1 Waves

Astronomical instrumentation, methods and techniques

03.01.2 Atmospheric effects 03.02.1 Balloons 03.09.1 Instrumentation: detectors 03.09.2 Instrumentation: interferometers 03.09.3 Instrumentation: miscellaneous 03.09.4 Instrumentation: photometers 03.09.5 Instrumentation: polarimeters 03.09.6 Instrumentation: spectrographs 03.13.1 Methods: analytical 03.13.2 Methods: data analysis 03.13.7 Methods: laboratory 03.13.3 Methods: miscellaneous 03.13.4 Methods: numerical 03.13.5 Methods: observational 03.13.6 Methods: statistical 03.19.1 Site testing 03.19.2 Space vehicles 03.20.1 Techniques: image processing 03.20.2 Techniques: interferometric 03.20.3 Techniques: miscellaneous 03.20.4 Techniques: photometric 03.20.5 Techniques: polarimetric 03.20.6 Techniques: radar astronomy 03.20.7 Techniques: radial velocities 03.20.8 Techniques: spectroscopic 03.20.9 Telescopes

Astronomical data bases

04.01.1 Astronomical data bases: miscellaneous 04.01.2 Atlases 04.03.1 Catalogs 04.19.1 Surveys Astrometry and celestial mechanics

05.01.1 Astrometry 05.03.1 Celestial mechanics, stellar dynamics 05.05.1 Eclipses 05.05.2 Ephemerides 05.15.1 Occultations 05.18.1 Reference systems 05.20.1 Time

The Sun

06.01.1 Sun: abundances 06.01.2 Sun: activity 06.01.3 Sun: atmosphere 06.03.1 Sun: chromosphere 06.03.2 Sun: corona 06.05.1 Sun: evolution 06.06.1 Sun: faculae, plages 06.06.2 Sun: filaments 06.06.3 Sun: flares 06.06.4 Sun: fundamental parameters 06.07.1 Sun: general 06.07.2 Sun: granulation 06.09.2 Sun: infrared 06.09.1 Sun: interior 06.13.1 Sun: magnetic fields 06.15.1 Sun: oscillations 06.16.1 Sun: particle emission 06.16.2 Sun: photosphere 06.16.3 Sun: prominences 06.18.1 Sun: radio radiation 06.18.2 Sun: rotation 06.19.1 (Sun:) solar-terrestrial relations 06.19.2 (Sun:) solar wind 06.19.3 (Sun:) sunspots 06.20.1 Sun: transition region 06.21.1 Sun: UV radiation 06.24.1 Sun: X-rays, gamma rays

Solar system

07.03.1 Comets: general 07.03.2 Comets: individual: … 07.05.1 Earth 07.09.1 Inteplanetary medium 07.13.1 Meteors, meteoroids 07.13.2 Minor planets, asteroids 07.13.3 Moon 07.16.1 Planets and satellites: general 07.16.2 Planets and satellites:

individual: … (alphabetic order)

07.19.1 Solar system: formation 07.19.2 Solar system: general Stars

08.01.1 Stars: abundances 08.01.2 Stars: activity 08.16.4 Stars: AGB and post-AGB 08.01.3 Stars: atmospheres 08.02.1 (Stars:) binaries (including multiple): close 08.02.2 (Stars:) binaries: eclipsing 08.02.3 (Stars:) binaries: general 08.02.4 (Stars:) binaries: spectroscopic 08.02.5 (Stars:) binaries: symbiotic 08.02.6 (Stars:) binaries: visual 08.02.7 (Stars:) blue stragglers 08.03.1 Stars: carbon 08.03.2 Stars: chemically peculiar 08.03.3 Stars: chromospheres 08.03.4 (Stars:) circumstellar matter 08.03.5 Stars: coronae 08.04.1 Stars: distances 08.05.1 Stars: early-type 08.05.2 Stars: emission-line, Be 08.05.3 Stars: evolution 08.06.1 Stars: flare 08.06.2 Stars: formation 08.06.3 Stars: fundamental parameters

(classification, colors, luminosities, masses, radii, temperatures, etc.)

08.07.1 Stars: general 08.08.1 (Stars:) Hertzsprung–Russel (HR) diagram 08.08.2 Stars: horizontal-branch 08.09.1 Stars: imaging 08.09.2 Stars: individual: … 08.09.3 Stars: interiors 08.11.1 Stars: kinematics 08.12.1 Stars: late-type 08.12.2 Stars: low-mass, brown dwarfs 08.12.3 Stars: luminosity function, mass function 08.13.1 Stars: magnetic fields 08.13.2 Stars: mass-loss 08.14.1 Stars: neutron 08.14.2 (Stars:) novae, cataclysmic variables 08.15.1 Stars: oscillations (including pulsations) 08.16.1 Stars: peculiar (except chemically peculiar) 08.16.2 (Stars:) planetary systems 08.16.3 Stars: Population II 08.16.5 Stars: pre-main sequence 08.16.6 (Stars:) pulsars: general 08.16.7 (Stars:) pulsars: individual: … 08.18.1 Stars: rotation 08.19.1 Stars: statistics 08.19.2 (Stars:) subdwarfs 08.19.3 (Stars:) supergiants 08.19.4 (Stars:) supernovae: general 08.19.5 (Stars:) supernovae: individual: … 08.19.6 (Stars:) starspots 08.22.1 (Stars: variables:) Cepheids 08.22.2 (Stars: variables:) δ Scu 08.22.3 Stars: variables: other 08.23.1 (Stars:) white dwarfs 08.23.2 Stars: Wolf-Rayet

Interstellar medium (ISM), nebulae

09.01.1 ISM: abundances 09.01.2 ISM: atoms 09.02.1 ISM: bubbles 09.03.1 ISM: clouds 09.03.2 (ISM:) cosmic rays 09.04.1 (ISM:) dust, extinction 09.07.1 ISM: general 09.08.1 (ISM:) H ii regions 09.09.1 ISM: individual objects: …

(except planetary nebulae)

09.10.1 ISM: jets and outflows 09.11.1 ISM: kinematics and dynamics 09.13.1 ISM: magnetic fields 09.13.2 ISM: molecules 09.16.1 (ISM:) planetary nebulae: general 09.16.2 (ISM:) planetary nebulae: individual: … 09.18.1 (ISM:) reflection nebulae 09.19.1 ISM: structure 09.19.2 ISM: supernova remnants

The Galaxy

10.01.1 Galaxy: abundances 10.03.1 Galaxy: center 10.05.1 Galaxy: evolution 10.06.1 Galaxy: formation 10.06.2 Galaxy: fundamental parameters 10.07.1 Galaxy: general 10.07.2 (Galaxy:) globular clusters: general 10.07.3 (Galaxy:) globular clusters: individual: … 10.08.1 Galaxy: halo 10.11.1 Galaxy: kinematics and dynamics 10.15.1 (Galaxy:) open clusters and associations: general 10.15.2 (Galaxy:) open clusters and associations:

individual: …

10.19.1 (Galaxy:) solar neighbourhood 10.19.2 Galaxy: stellar content 10.19.3 Galaxy: structure

Galaxies

11.01.1 Galaxies: abundances 11.01.2 Galaxies: active 11.02.1 (Galaxies:) BL Lacertae objects: general 11.02.2 (Galaxies:) BL Lacertae objects: individual:

…

11.03.1 Galaxies: clusters: general 11.03.4 Galaxies: clusters: individual: … 11.03.2 Galaxies: compact 11.03.3 (Galaxies:) cooling flows 11.04.1 Galaxies: distances and redshifts 11.05.1 Galaxies: elliptical and lenticular, cD 11.05.2 Galaxies: evolution 11.06.1 Galaxies: formation 11.06.2 Galaxies: fundamental parameters

(classification, colors, luminosities, masses, radii, etc.)

11.07.1 Galaxies: general 11.08.1 Galaxies: halos 11.09.1 Galaxies: individual: … 11.09.2 Galaxies: interactions 11.09.3 (Galaxies:) intergalactic medium 11.09.4 Galaxies: ISM 11.09.5 Galaxies: irregular 11.10.1 Galaxies: jets 11.11.1 Galaxies: kinematics and dynamics 11.12.1 (Galaxies:) Local Group 11.12.2 Galaxies: luminosity function, mass function 11.13.1 (Galaxies:) Magellanic Clouds 11.13.2 Galaxies: magnetic fields 11.14.1 Galaxies: nuclei 11.16.1 Galaxies: photometry 11.16.2 Galaxies: peculiar 11.17.1 (Galaxies:) quasars: absorption lines 11.17.2 (Galaxies:) quasars: emission lines 11.17.3 (Galaxies:) quasars: general 11.17.4 (Galaxies:) quasars: individual: … 11.19.1 Galaxies: Seyfert 11.19.2 Galaxies: spiral 11.19.3 Galaxies: starbust 11.19.4 Galaxies: star clusters 11.19.7 Galaxies: statistics 11.19.5 Galaxies: stellar content 11.19.6 Galaxies: structure

Cosmology

12.03.1 (Cosmology:) cosmic microwave background 12.03.2 Cosmology: miscellaneous 12.03.3 Cosmology: observations 12.03.4 Cosmology: theory 12.04.1 (Cosmology:) dark matter 12.04.2 (Cosmology:) diffuse radiation 12.04.3 (Cosmology:) distance scale 12.05.1 (Cosmology:) early Universe 12.07.1 (Cosmology:) gravitational lensing 12.12.1 (Cosmology:) large-scale structure of Universe

Sources as a function of wavelength

13.07.1 Gamma rays: bursts 13.07.2 Gamma rays: observations 13.07.3 Gamma rays: theory 13.09.1 Infrared: galaxies 13.09.2 Infrared: general 13.09.3 Infrared: ISM: continuum 13.09.4 Infrared: ISM: lines and bands 13.09.5 Infrared: solar system 13.09.6 Infrared: stars 13.18.1 Radio continuum: galaxies 13.18.2 Radio continuum: general 13.18.3 Radio continuum: ISM 13.18.4 Radio continuum: solar system 13.18.5 Radio continuum: stars 13.19.1 Radio lines: galaxies 13.19.2 Radio lines: general 13.19.3 Radio lines: ISM 13.19.4 Radio lines: solar system 13.19.5 Radio lines: stars 13.21.1 Ultraviolet: galaxies 13.21.2 Ultraviolet: general 13.21.3 Ultraviolet: ISM 13.21.4 Ultraviolet: solar system 13.21.5 Ultraviolet: stars 13.25.1 X-rays: bursts 13.25.2 X-rays: galaxies 13.25.3 X-rays: general 13.25.4 X-rays: ISM 13.25.5 X-rays: stars